OpenStudy (anonymous):

CHEMICAL EQUILIBRIUM!! Consider the following reversible reaction. In a 3.00-liter container, the following amounts are found in equilibrium at 400°C: 0.0420 mole N2, 0.516 mole H2, and 0.0357 mole NH3. Evaluate Kc . N2(g) + 3H2(g) <---> 2NH3(g) a. 0.202 b. 1.99 c. 16.0 d. 4.94 e. 0.503 Answer is B I understand the setup but what do I do with the 3.00L and 400 degrees C? How could I get the answer b??

OpenStudy (frostbite):

The equilibrium constant at 400 degrees C for the reaction: \[\large Kc _{T=400}=\frac{ \left[ NH _{3}(g) \right]^{2} }{ \left[ N _{2}(g) \right] \left[ H _{2}(g) \right]^{3} }\] The concentrations are defined as: \[\large C=\frac{ n }{ V }\]

OpenStudy (frostbite):

Substitute and we get the following: \[\LARGE Kc _{T=400}=\frac{ \left( \frac{ n(NH _{3}(g)) }{ 3.00l } \right)^{2} }{ \frac{ n(N _{2}(g)) }{ 3.00l } \left( \frac{ n(H _{2}(g)) }{ 3.00l } \right)^{3}}\]

OpenStudy (frostbite):

For evaluation see attachment.

OpenStudy (frostbite):

All the measurements have been done at 400 degrees C therefor the equilibrium constant is also at 400 degrees C.

OpenStudy (anonymous):

When calculating the answer the way it was set up in your attachment the answer comes out to 1.987 x 10 ^-10. The answer you gave was just the 1.987 which would correspond to the supposed correct answer as it is given in option b which is 1.99. I'm confused how I got the answer but with a power of -10?

OpenStudy (frostbite):

Try look if your brackets are set correctly else try calculate it in parts. Or you can take a screenshot if using a calculator on your PC and let me go through the calculation.